please assist .


There are 300 students at a college that prepare students for foreign universities. The college offers three foreign languages, English, French and German. The director of studies at the center provided the following information concerning the environment of students in these three languages. The information indicated that 100 students enrolled in English, 80 in French, 50 in german, 16 in English and French, 10 in English and German, 8 in French and German while 1 student enrolled in all the 3 languages. A student is selected at random from the college. Find the probability that the student is enrolled in
a) 3 languages
b) At least 2 languages
c) 1 language only
d) None of the languages

Answer by stanbon(48568)

You have to draw the Venn diagram and work from the
intersection of the three circles outward
E and F and G = 1
E and F = 16
E and G =10
G and F = 8
E only =75
F only =57
G only = 3
Total with a language = 230
No language = 70
---------------
a) P(3 languages)=1/300
b) P(At least 2 languages)=(32+1)/300=33/300
c) P(1 language only)=(75+57+33)/300
d) P(None of the languages)=70/300
------------
Cheers,
Stan H.